OpenStudy (anonymous):

prove the identity: (cosx/1-sinx)+(cosx/1+sinx) =2secx

OpenStudy (anonymous):

find the common denominator

OpenStudy (anonymous):

1-sin^2(x) ( DIFFERENCE OF SQUARES ) which turns in cos^2(x)

OpenStudy (anonymous):

LHS should be 2cos(x)/cos^2(x)

OpenStudy (anonymous):

first combine the fractions by getting a common denominator and adding \[(cosx(1+sinx) + cosx(1-sinx))\div (1-sinx)(1+sinx)\] now factor out the Cosx from the top and expand the bottom to get: \[(cosx (1+ sinx + 1 - sinx)) \div (1 - Sin^2x)\] simplifying we get: \[cosx (2) \div (1-\sin^2x)\] Using the Pathagorean Identity, we know that \[1-\sin^2 x = \cos^2x\] so we can rewrite it as \[(2)cosx \div \cos^2x\] the top and bottom cancel to become (2)/cosx = 2x (1/cosx) since secx = 1/cosx it becomes 2secx

OpenStudy (anonymous):

sorry the last part shoudl say: (2)/cosx = 2(1/cosx)

OpenStudy (anonymous):

Hermeezey you have to give the student hints not solutions ...

OpenStudy (anonymous):

hmm ok i just got stuck on the top part

OpenStudy (anonymous):

which part do you mean?

OpenStudy (anonymous):

what happend to the sinx's for the numerator?

OpenStudy (anonymous):

Lol I see, i kinda skimmed that part i guess. because (1+ sinx + 1 - sinx) can be reorganized as (1 + 1 + sinx - sinx) and we have a +sinx and -sinx, they just cancel out, so you are left with a (1+1).

OpenStudy (anonymous):

oh i see. ok i also so dont get how 2cosx became 2/cosx...

OpenStudy (anonymous):

ok i understand now! thanks